OpenStudy (anonymous):

-11=(1/6) X + 9.......solve for x please.

OpenStudy (diyadiya):

Subtract 9 from both sides -20 = (1/6) x Now multiply both sides by 6

OpenStudy (diyadiya):

Can you do it?

OpenStudy (anonymous):

-11-9=1/6x -20x6=x x=-120